OpenStudy (anonymous):

what is the equation of the tangent line to y= e^4x + tan-1 (x) at x=0

OpenStudy (anonymous):

tangent line = derivative at the point for slope, then just use the point. So derivative of this is 4e^(4x)+1/(x^2+1). At x=0 it is 4+1=5. The point is then (0,0). So use y-y1=m(x-x1) to get y-0=5(x-0), or y=5.

myininaya (myininaya):

the point is (0,1) that we are trying to find the tangent line

OpenStudy (anonymous):

Point is actually (0,1), so line is y-1=5(x-0)

myininaya (myininaya):

gj mark

OpenStudy (anonymous):

I didn't know what the y point was how do you find it?

OpenStudy (anonymous):

Plug in x=0 into the original equation

OpenStudy (anonymous):

So is e^4x 1 then? Thats what I'm confused about

myininaya (myininaya):

e^(0)=1

myininaya (myininaya):

just like 5^(0)=1 and 2^(0)=1

OpenStudy (anonymous):

Ooh ok thanks

myininaya (myininaya):

if tan^(-1)(0)=s then tan(s)=0 what value of s betweeen -pi/2 and pi/2 makes this expressions true

OpenStudy (anonymous):

Ok got it now... thanks so much

myininaya (myininaya):

np

myininaya (myininaya):

when i said expression i mean equation lol

OpenStudy (anonymous):

yea gotcha